smple of s mconded dn the United States follows a normal distribution. A random sample of 64 recorded deaths in the United of 71 8 years with a standard deviation of S.9 years. (20pes) 6 Assume that the life span in follows a States during the past year showed an average life span (a) (10pts) Construct intervail of the mcan life span narrower or wider? Explain your reasoning. bi (00pe) Does ycars? Use a 005 level of significance and follow the hypothesis test procedures a95% confidence interval of the mean life span, would a 90% confidence the data collected seem to indicate that the mean life span today is greater than 70
Answer #1

Solution: Sample size = n = 64 Sample mean =x = 71.8 Sample Standard Deviation =s 8.9 Degrees of freedom =df = n-1 64-1 = 63Part-B: Hai μ 70 H1: μ> 70 Test Statistic: t= (718-70)/(89/ 「64) t = 1.618 From t-distribution table, area under the curve to